In the Bohr model of the hydrogen atom, the allowed orbits of the electron (labeled n = 1, 2, 3, …) have angular momentum , orbital radii , and energies . In these expressions me is the mass of the electron.

In an exotic atom the electron is replaced by a different subatomic particle that has the same charge as an electron but a different mass. Two examples that have been studied are muonic hydrogen, in which the electron is replaced by a muon of mass , and kaonic hydrogen, in which the electron is replaced by a kaon of mass .

(a) Rank the radii of the lowest-energy orbits in (i) ordinary hydrogen, (ii) muonic hydrogen, and (iii) kaonic hydrogen, from largest radius to smallest radius. If any two of these orbits have the same radius, state this. Explain how you made your ranking.

(d) Rank the speeds of the particles orbiting the nucleus in the lowest-energy orbit of (i) ordinary hydrogen, (ii) muonic hydrogen, and (iii) kaonic hydrogen, from fastest to slowest. If the speed is the same in any two of these situations, state this. Explain how you made your ranking.
Answer #1

1) We know, bohr's atomic model states that the orbits of electron( or muon/kaon) must have quantised angular momentum.

i.e. mvr=\frac{nh}{2\pi }               

vr=\frac{nh}{2\pi m }                 ----------(1)

Also, the electrostatic force between the proton and electron provides the centripetal force for rotation. so,

\frac{q^{2}}{4\pi \epsilon _{0}r^{2}} = mv^{2}/r => v^{2}r^{2}= \frac{q^{2}r}{4\pi \epsilon _{0}m}

Substituting LHS from (1), we get

\frac{n^{2}h^{2}\epsilon_{0}}{q^{2}\pi m} = r             -------(2)

so, radius of orbit is inversely proportional to mass or particle orbiting.

Now, Mass of Kaon >Mass of muon >mass of electron

So, radius(kaon)<radius(muon)<radius(electron)

2) Now, from (1),

v=\frac{nh}{2\pi m r}

and using r from (2), we get

v=\frac{q^{2}}{2nh\epsilon _{0}}

So, velocity of orbiting particle is independent of the mass , and hence will be same for all three , electron kaon and muon.